Nano Network Inc.’s financial data are listed in Table below:

Compute the financial ratios during the fiscal year 2009 and interpret the company’s financial health. The closing stock price was $ 5.67 on Dec. 31, 2009, and the outstanding shares are 90,340,000. Note the balance sheet and income statement entries in this problem are not complete. Only relevant entries are listed. Do not attempt to add individual entries to confirm either current assets or current liabilities. Some ratios may not be defined due to lack of relevant financial data, just indicate “cannot determine it/not defined”.

  1. Debt ratio
  2. Times-interest-earned ratio
  3. Current ratio
  4. Quick ratio
  5. Inventory turnover ratio
  6. Day’s sales outstanding
  7. Total assets turnover
  8. Profit margin on sales
  9. Return on total assets
  10. Return on common equity
  11. Price-to-earnings ratio
  12. Book value per share

TABLE P2.3Balance Sheet for Nano Networks Inc.

Dec. 2009

U.S. $ (000)

(Year)

Dec. 2008

U.S. $ (000)

(Year)

Balance Sheet

Summary

Cash

158,043

20,098

Securities

285,116

0

Receivables

24,582

8,056

Inventory

0

0

Current assets

377,833

28,834

Property and equipment, net

20,588

10,569

Depreciation

8,172

2,867

Total assets

513,378

36,671

Current liabilities

55,663

14,402

Bonds

0

0

Preferred stock

0

0

Total common stockholder’s equity

457,715

17,065

Total liabilities and equity

513,378

36,671

Income Statement

Summary

Total revenues

102,606

3,807

Cost of sales

45,272

4,416

Other expenses

71,954

31,661

Interest income

8,011

1,301

Income pre-tax

-6,609

-69

Income tax

2,425

2

Income continuing

-9,034

-30,971

Net income

-9,034

-30,971

EPS primary

-$0.1

-$0.80
Answer #1

As per rules I am answering the first 4 subparts of the question

  1. Debt ratio = Debt/ Totalassets

= ( 513378-457715)/ 513378

=10.84%

Debt is very low. In fact the company can think of further financing through debt.

  1. Times-interest-earned ratio = EBIT/ Interest

= (102606-45272-71954)/ 8011

=- 1.382

The EBIT is negative and not enough to cover the interest payments.

  1. Current ratio=Current assets/ current liabilities

= 377833/55663

=6.79

This ratio is very high to cover the liabilities showing good liquidity of the business.

  1. Quick ratio = (Current assets – Inventory) / Current liabilities

= (377833-0) /55663

=6.79

This ratio is very high to cover the liabilities showing good liquidity of the business.